# Prediction-Feedback DETR for Temporal Action Detection

### v2

@@ Introduction
TriDet (Shi et al. 2023) I3D $\boldsymbol{83.6}$ $\boldsymbol{80.1}$ $\boldsymbol{72.9}$ $\boldsymbol{62.4}$ $\boldsymbol{47.4}$ $\boldsymbol{69.3}$ TSP $54.70$ $\boldsymbol{38.00}$ $8.40$ $\boldsymbol{36.80}$
@@
Pred-DETR (Ours) I3D $\boldsymbol{80.0}$ $\boldsymbol{73.5}$ $\boldsymbol{64.6}$ $\boldsymbol{52.3}$ $\boldsymbol{37.6}$ $\boldsymbol{61.6}$ I3D $\boldsymbol{58.38}$ $\boldsymbol{39.14}$ $9.92$ $\boldsymbol{38.62}$
Table 1: The comparison results with the state-of-the-art on THUMOS14 and ActivityNet-v1.3. ‘Feat.’ indicates the backbone features including I3D, TSN and TSP (R(2+1)D).
@@
Table 2: The comparison results on the HACS dataset.$\dagger$’ indicates our reproduced version, and SF means the features from the SlowFast backbone.
@@
Table 3: The comparison results on the FineAction dataset.$\dagger$’ indicates our reproduced version. VM2 indicates VideoMAEv2-g features.
@@
In the table, our model remarkably outperforms all DETR-based models on all benchmarks. It demonstrates that when the attention collapse is relieved, the original-DETR architecture can be comparable or superior to the Deformable-DETR architecture in TAD, aligned with the observation in object detection (Lin et al. 2023). More interestingly, Pred-DETR best performs on ActivityNet-v1.3 including the non-DETR methods. DETR-based methods tend to produce better results on ActivityNet and HACS than those on THUMOS14 and FineAction. This could be because ActivityNet and HACS mostly contain long actions while THUMOS14 and FineAction includes many short instances. Predicting short actions precisely requires high temporal resolution while DETR does not handle yet such a long sequence due to the query-based architecture. Nevertheless, recent DETR models including ours show superior performances over the non-DETR models which handle short-length sequences except for ActionFormer and TriDet.
@@
Figure 4: Diversity of attention maps. The qualitative measure of the diversity for the cross- and self-attention maps for the test samples of ActivityNet-v1.3.
@@
Diversity. To clearly validate the effect of the prediction-feedback for the attention collapse, we measure the diversity of cross- and self-attention maps according to (Dong, Cordonnier, and Loukas 2021; Kim, Lee, and Heo 2023). The diversity $d(A)$ for the attention map $A$ is the measure of the closeness between the attention map and a rank-1 matrix as defined as below:
@@
Fig. 4 shows the diversity on each layer of the encoder and decoder for the baseline DETR and Pred-DETR. The diversity is measured on the test set on ActivityNet-v1.3 averaged over all test samples. As the model depth gets deeper, the diversity of the baseline decreases close to 00. However, the diversity of Pred-DETR does not fall down and even increases. From these results, our prediction-feedback effectively relieves the collapse problem.

### v3

@@ Introduction
TriDet (Shi et al. 2023) I3D $83.6$ $\boldsymbol{80.1}$ $\boldsymbol{72.9}$ $\boldsymbol{62.4}$ $47.4$ $\boldsymbol{69.3}$ TSP $54.70$ $38.00$ $8.40$ $36.80$
DyFaDet (Yang et al. 2024) I3D $\boldsymbol{84.0}$ $\boldsymbol{80.1}$ $72.7$ $61.1$ $\boldsymbol{47.9}$ $69.2$ TSP $\boldsymbol{58.10}$ $\boldsymbol{39.60}$ $8.40$ $\boldsymbol{38.50}^{*}$
@@
Pred-DETR (Ours) I3D $80.0$ $73.5$ $64.6$ $52.3$ $37.6$ $61.6$ I3D $54.17$ $36.43$ $9.53$ $36.00$
Pred-DETR (Ours) VM2 $\boldsymbol{84.1}$ $\boldsymbol{80.0}$ $\boldsymbol{72.2}$ $\boldsymbol{60.4}$ $\boldsymbol{45.8}$ $\boldsymbol{68.5}$ I3D $\boldsymbol{58.38}$ $\boldsymbol{39.14}$ $9.92$ $\boldsymbol{38.62}^{*}$
Table 1: The comparison results with the state-of-the-art on THUMOS14 and ActivityNet-v1.3. ‘Feat.’ indicates the backbone features including I3D, TSN, TSP (R(2+1)D) and VM2. ‘*’ means using UniFormer-v2 classification on ActivityNet-v1.3.
@@
Table 2: The comparison results on the HACS dataset.$\dagger$’ indicates our reproduced version.
@@
Table 3: The comparison results on the FineAction dataset.$\dagger$’ indicates our reproduced version.
@@
In the table, our model remarkably outperforms all DETR-based models on all benchmarks. It demonstrates that when the attention collapse is relieved, the original-DETR architecture can be comparable or superior to the Deformable-DETR architecture in TAD, aligned with the observation in object detection (Lin et al. 2023). More interestingly, Pred-DETR best performs on ActivityNet-v1.3 including the non-DETR methods. DETR-based methods tend to produce better results on ActivityNet and HACS than those on THUMOS14 and FineAction. This could be because ActivityNet and HACS mostly contain long actions while THUMOS14 and FineAction include many short instances. Predicting short actions precisely requires high temporal resolution while DETR does not handle yet such a long sequence due to the query-based architecture. Nevertheless, recent DETR models including ours show superior performances over the non-DETR models which handle short-length sequences except for ActionFormer and TriDet.
@@
Figure 4: Diversity of attention maps. Diversity for cross- and self-attention for test samples of ActivityNet-v1.3.
@@
Diversity. To clearly validate the effect of the feedback for the collapse, we measure the diversity of cross- and self-attention maps according to (Dong, Cordonnier, and Loukas 2021; Kim, Lee, and Heo 2023). The diversity $d(A)$ for the attention map $A$ is the measure of the closeness between the attention map and a rank-1 matrix as below:
@@
Fig. 4 shows the diversity on each layer of the encoder and decoder for the baseline DETR and Pred-DETR. The diversity is measured on the test set on ActivityNet-v1.3 averaged over all test samples. As the model depth gets deeper, the diversity of the baseline decreases close to 00. However, the diversity of Pred-DETR does not fall down and even increases. From this, Pred-DETR effectively relieves the collapse.
@@ Conclusion
Acknowledgements
This work was supported in part by MSIT/IITP (No. 2022-0-00680, 2020-0-01821, 2019-0-00421, RS-2024-00459618, RS-2024- 00360227, RS-2024-00437102, RS-2024-00437633), and MSIT/NRF (No. RS-2024-00357729).